Profit Margin, Investment Turnover, and Return on Investment The condensed income statement for the Consumer Products...

  1. Profit Margin, Investment Turnover, and Return on Investment

    The condensed income statement for the Consumer Products Division of Tri-State Industries Inc. is as follows (assuming no support department allocations):

    Sales $996,000
    Cost of goods sold (448,200)
    Gross profit $547,800
    Administrative expenses (199,200)
    Operating income $348,600

    The manager of the Consumer Products Division is considering ways to increase the return on investment.

    a. Using the DuPont formula for return on investment, determine the profit margin, investment turnover, and return on investment of the Consumer Products Division, assuming that $1,660,000 of assets have been invested in the Consumer Products Division. Round the investment turnover to one decimal place.

    Profit margin %
    Investment turnover
    Return on investment %

    b. If expenses could be reduced by $49,800 without decreasing sales, what would be the impact on the profit margin, investment turnover, and return on investment for the Consumer Products Division? Round the investment turnover to one decimal place.

    Profit margin %
    Investment turnover
    Return on investment %

Check My Work
Answer #1
Please give positive ratings so I can keep answering. It would help me a lot. Please comment if you have any query. Thanks!
Tri-State Industries Inc.
Answer a Amount $ Note
Net Sales      996,000.00 A
Operating income      348,600.00 B
Total Assets 1,660,000.00 C
As per Du Pont, Return on investment= Profit margin*Investment turnover.
Profit margin 35.00% D=B/A
Investment turnover                  0.60 E=A/C
Return on investment 21.00% F=D*E
Answer b Amount $ Note
Net Sales      996,000.00 A
Operating income      398,400.00 G= B+49,800. If expenses will decrease then operating income will increase.
Total Assets 1,660,000.00 C
As per Du Pont, Return on investment= Profit margin*Investment turnover.
Profit margin 40.00% H=G/A
Investment turnover                  0.60 I=A/C
Return on investment 24.00% J=H*I